The distribution of grade point averages​ (GPAs) for medical school applicants of a certain year were approximately​ Normal, with a mean of 3.55 and a standard deviation of 0.33. Suppose a medical school will only consider candidates with GPAs in the top 20​% of the applicant pool. An applicant has a GPA of 3.84. Does this GPA fall in the top 20​% of the applicant​ pool?

Select the correct choice below and fill in the answer box to complete your choice.

​(Type an integer or decimal rounded to two decimal places as​needed.)

A. Yes. The cutoff for the top 20​% is a GPA of blank

B.No. The cutoff for the top 20​% is a GPA of blank


1
Expert's answer
2021-07-22T18:23:26-0400

Let X the random variable that represents the grade point averages (GPAs) for medical school applicants of a certain year of a population, and for this case we know the distribution for X is given by:

X~N(3.55,0.33)

"\mu=3.55 \\

\sigma= 0.33"

The best way to solve this problem is using the normal standard distribution and the z score given by:

"Z = \frac{x - \mu}{\sigma}"

We want to find a value a, such that we satisfy this condition:

"P(X<a)=0.80 \\

P(Z< \frac{a- \mu}{\sigma}) = 0.80 \\

\frac{a- 3.55}{0.33} = 0.842 \\

a = 3.55+0.33 \times 0.842 = 3.827"

So the score that separates the bottom 80 % of data from the top 20 % is 3.827. Since the value obtained by the applicant is 3.84 >3.827 falls on the top 20 %.

Answer: A. Yes. The cutoff for the top 20​% is a GPA of blank